# Meta SecAlign: A Secure Foundation LLM Against Prompt Injection Attacks

### v2

@@ V-E Meta SecAlign has trivial utility drop to enable prompt injection security
TABLE VII: Utility When using An LLM with or without prompt injection defense (prompt-data channel separation).
Model PI Defense GPT Gemini-Flash
Meta-SecAlign-70B -4o-mini -4o -5 (high) -2.0 -2.5
AlpacaEval2 Utility ($\uparrow$) No 44.2% 52.2% 62.4% 70.1% 51.3% 69.0%
AlpacaEval2 Utility ($\uparrow$) Yes 44.7% 44.7% 56.4% 68.7% 38.8% 44.6%
Difference +0.5% -7.5% -6.0% -1.4% -12.5% -22.4%
SEP Utility ($\uparrow$) No 62.1% 67.9% 76.0% 76.0% 64.0% 68.7%
SEP Utility ($\uparrow$) Yes 60.4% 62.1% 62.5% 76.8% 38.2% 49.5%
Difference -1.7% -5.8% -13.5% +0.8% -25.8% -19.2%
A prompt injection defense is implemented by using separate channels to accept the prompt and the data [chen2024struq]. Here, we study the utility drop for this system’s channel separation. That is, will the utility scores increase if we discard our security goal, i.e., by putting both the prompt text and the data text into the prompt input channel (within the user message role)?
The answer is no for Meta-SecAlign-70B and GPT-5, meaning that the model’s full utility has already been unlocked when PI security is turned on during the prompt-data channel separation, see Table˜VII. However, this is not the case for GPT-4o-mini, GPT-4o, Gemini-2.0-Flash, and Gemini-2.5-Flash, where putting prompt and data in separate message types for PI security noticeably hurts the utility.
Meta SecAlign achieves this good property (no utility drop for security) by using the new input message role for untrusted texts in a free-form manner. For PI security, any texts that were put in the user role can now be directly put within input delimiters if they are untrusted, see Section˜III-A.
For all tested commercial LLMs, the prompt-data separation is established between the user role and the tool role [wallace2024hierarchy, shi2025lessons], which only treats the tool return as the untrusted data. In our evaluations, we have to create a dummy tool that returns the data texts, and asks the model to call it so that the untrusted data can be processed securely as designed. We cannot know if this design correlates with "the utility drop for PI security" as those models are proprietary, but we hypothesize that the system designs for tool-calling may lead to utility drop in instruction-following tasks as in Section˜III-A. GPT-5 may address this issue by routing proper sub-models to solve corresponding tasks, so we do not observe any utility difference when PI security is turned on or not.
V-F Stronger LLMs are more vulnerable to PIs if left undefended
As LLMs become more capable at instruction following, any injected instructions tend to be easily identified and thus followed, leading to vulnerability to PIs. We test three undefended Llama 3 LLMs, sorted by capability in ascending order, see Figure˜4. Without defense, stronger LLMs suffer from consistently higher ASRs as hypothesized. Fortunately, Figure˜4 (bottom) shows that after SecAlign++, all three LLMs can reach a similarly good level of security. This trend gives us hope that, as frontier LLMs continue to improve in capabilities, it is still possible to effectively secure them against PI attacks. But defenders’ moves are more and more pressing; otherwise, the undefended model becomes a perfect target for attackers.
@@ V-F Stronger LLMs are more vulnerable to PIs if left undefended
Figure 4: Security (attack success rate $\downarrow$) of LLMs with different instruction-following capabilities (Llama-3.1-8B-Instruct < Llama-3.1-70B-Instruct < Llama-3.3-70B-Instruct). Stronger LLMs are more vulnerable to PI attacks when undefended (top), but could be fine-tuned to a similar level of robustness (bottom). Detailed numbers are present in Table˜XI.
VI Discussions
VI-A Conclusions
@@
Model-level defenses are currently most effective in mitigating prompt injections, the top-1 threat to LLM agents. Prior model-level defenses are either tested in toy 8B LLMs in academic papers or deployed in closed-source industry APIs (except the recent open-weight-only GPT-OSS [agarwal2025gpt] after our release). Our work bridges this gap by fully open-sourcing a commercial-grade robust foundation LLM, Meta-SecAlign-70B, for securing agentic applications. Our training recipe, SecAlign++, for the first time, achieves significant security with no noticeable utility drop in the most comprehensive evaluations to date. Interestingly, this security generalizes to diverse downstream tasks unseen in SecAlign++, especially in agentic workflows where prompt injection is the major threat. We hope our work can inspire and accelerate the co-development of PI attacks and defenses in the community.
VI-B Future Directions
Defending against stronger prompt injections. Our models are robust in major PI benchmarks with optimization-free attacks, where the attackers have limited computation and knowledge about the target model. However, the PI problem is not solved - recent optimization-based attacks [nasr2025attacker, wen2025rl, pandya2025may] have shown near 100% attack success rates on Meta SecAlign by iteratively optimizing the injection. Future work is needed to mitigate those strong adaptive attacks.
@@ VI-B Future Directions
Online reinforcement learning for securing reasoning LLMs. Our training recipe relies on offline preference optimization to build PI security policy into the non-reasoning model. Recently, online reinforcement learning, such as GRPO [guo2025deepseek], has unlocked better model reasoning using limited data [guo2025deepseek]. Applying it to defending against PI attacks, e.g., by designing PI security rewards, may enjoy similar benefits, with the major challenge lying in utility preservation. Reasoning LLMs, with stronger instruction-following abilities, are more vulnerable to PIs [evtimov2025wasp], and the only work of securing them against PIs is [wu2025thinkingcontrol] to our knowledge.
@@
Defending against visual prompt injections. In contrast to textual attacks, the image modality is continuous in nature, and vision models have traditionally been more vulnerable to adversarial manipulation [carlini2019evaluating]. Research towards this problem is especially relevant for agentic web navigation, as state-of-the-art web agents such as OpenAI Operator [openai_operator_system_card], Claude Computer Use [anthropic_claude_computer_use], and Google DeepMind’s Project Mariner [deepmind_project_mariner] are typically powered by multi-modal models.
VI-C Limitations
We focus on defending against (indirect) PIs, where the user is benign, but the environment is malicious, as in agents. Thus, our work cannot prevent jailbreak [zou2023universal], direct PI [mu2025closer], and other attacks. Direct PIs are also called system message following attacks where the user is malicious, trying to override the system message, so for defense, the user prompt should be conditionally followed only if it does not conflict with the system. Defenses by contextually identifying instructions to follow are mitigating direct PIs, while defenses by enhancing the system’s prompt-data separation are mitigating indirect PIs, where which input part is trusted or untrusted is clear without needing to be determined by additional contexts.
We only consider single-turn PIs, where the attack payload is put in one turn of the data message. For AgentDojo and WASP, there are multi-turn conversations, but only one turn of the data is injected. There is no PI benchmark where the attack is progressively posted, to the best of our knowledge.
Our training recipe has only been tested in the Llama 3 series, and it is uncertain how much hyper-parameter tuning it requires to make SecAlign++ generalize to other models, especially those with MoE architectures and reasoning abilities.
Even though randomized injection position mitigates the learning of a specific shortcut, we still observe another learned shortcut in SecAlign++, forcing whoever uses Meta SecAlign to always put the prompt part (user role) before the data part (input role). This is acceptable in practice as the trusted system developer controls how the LLM is used. But scientifically, it is unclear if the model recognizes message types instead of their orders.
@@ Acknowledgments
In the main paper, we test GPT-5 on its high reasoning level with the best utility. In Table˜VIII, we present results with other reasoning levels, which show similar security scores. In the main paper, we present an easy control of the utility-security trade-off by tuning LoRA $\alpha$ at test time, see Figure˜3. This trade-off is traditionally controlled by tuning the learning rate at training time, see Figure˜5. In the main paper, we put sandwich prompting (Please always remember that your task is: {instruction}) at the end of data in InjecAgent and AgentDojo. In Table˜IX, we present results without the sandwich defense, which also support Meta SecAlign’s advantage.
TABLE VIII: Security and utility evaluations on GPT-5 with different reasoning levels.
@@
MMLU-Pro 5-shot ($\uparrow$) - 80.6% 87.1%
GPQA Diamond 0-shot ($\uparrow$) - 67.3% 85.1%
AlpacaEval2 ($\uparrow$) 67.8% 63.7% 68.7%
@@
Figure 5: Tuning learning rate at training time can also control the utility(left)-security(right) trade-off for SecAlign++ on Llama-3.3-70B-Instruct.
TABLE IX: InjecAgent and AgentDojo results without sandwich prompting defense, which is added on those two benchmarks in the main paper.
Model Llama-3.3-70B-Instruct GPT Gemini-Flash
Undefended Meta-SecAlign -4o-mini -4o -5 (high) -2.0 -2.5
InjecAgent ASR ($\downarrow$) 86.0% 2.1% 7.7% 36.9% 0.6% 66.5% 3.5%
AgentDojo Utility ($\uparrow$) 62.9% 79.4% 70.1% 80.4% 83.5% 44.3% 58.8%
AgentDojo Utility w. Attack ($\uparrow$) 41.9% 77.1% 40.6% 38.8% 81.3% 37.3% 42.9%
AgentDojo ASR ($\downarrow$) 23.0% 2.3% 30.9% 43.2% 0.2% 12.4% 30.7%
TABLE X: Numbers in Figure˜3: Performance (%) on using different LoRA $\alpha$ in Meta-SecAlign-70B
@@
Knowledge MMLU 0-shot ($\uparrow$) 86.3% 86.5% 86.2% 86.1% 86.0% 85.9% 85.9% 85.8% 85.9%
@@
GPQA Diamond 0-shot ($\uparrow$) 50.0% 48.0% 49.0% 49.0% 47.0% 48.5% 49.0% 47.5% 48.0%
Instrustion Following AlpacaEval2 ($\uparrow$) 44.2% 44.4% 45.5% 44.8% 45.0% 44.1% 44.5% 44.6% 44.7%
@@
AlpacaFarm Adaptive ASR ($\downarrow$) 98.1% 97.6% 96.2% 87.0% 59.1% 34.6% 6.3% 1.0% 0.5%
@@
Agentic Workflows InjecAgent Base ASR ($\downarrow$) 23.5% 13.1% 6.6% 3.4% 1.0% 0.4% 0.2% 0.2% 0.2%
InjecAgent Enhanced ASR ($\downarrow$) 53.6% 39.8% 20.9% 9.5% 4.2% 3.2% 1.3% 0.9% 0.2%
@@
TABLE XI: Numbers in Figure˜4: security and utility evaluations on LLMs with different powerfulness.
Undefended Model Llama-3.1-8B-Instruct Llama-3.1-70B-Instruct Llama-3.3-70B-Instruct
@@
GPQA Diamond 0-shot ($\uparrow$) 31.3% 28.3% 44.4% 43.9% 50.0% 48.0%
AlpacaEval2 ($\uparrow$) 31.2% 31.0% 43.6% 43.9% 44.2% 44.7%
@@
AlpacaFarm Adaptive ASR ($\downarrow$) 89.4% 0.5% 94.7% 0% 98.1% 0.5%

### v3

@@ 4.6 Meta SecAlign has trivial utility drop to enable prompt injection security
Table 8: Utility when using an LLM with or without prompt injection defense (prompt-data channel separation).
Defense GPT Gemini
Ours 4o-mini 4o 5 2-Flash 2.5-Flash 3-Pro
AlpacaEval2 Utility ($\uparrow$) No 44.2% 52.2% 62.4% 70.1% 51.3% 69.0% 67.7%
AlpacaEval2 Utility ($\uparrow$) Yes 44.7% 44.7% 56.4% 68.7% 38.8% 44.6% 64.3%
Difference +0.5% -7.5% -6.0% -1.4% -12.5% -22.4% -3.4%
SEP Utility ($\uparrow$) No 62.1% 67.9% 76.0% 76.0% 64.0% 68.7% 76.1%
SEP Utility ($\uparrow$) Yes 60.4% 62.1% 62.5% 76.8% 38.2% 49.5% 70.1%
Difference -1.7% -5.8% -13.5% +0.8% -25.8% -19.2% -6.0%
@@
Figure 4: Security (attack success rate $\downarrow$) of LLMs with different instruction-following capabilities (Llama-3.1-8B-Instruct < Llama-3.1-70B-Instruct < Llama-3.3-70B-Instruct). Stronger LLMs are more vulnerable to PI attacks when undefended (left), but could be fine-tuned to a similar level of robustness (right). Detailed numbers are present in Table˜12.
A prompt injection defense is implemented by using separate channels to accept the prompt and the data [chen2024struq]. Here, we study the utility drop under this system’s channel separation. That is, would the utility scores increase if we discard our security goal, i.e., by putting both the prompt text and the data text into the prompt input channel (within the user message role)?
The answer is no for Meta-SecAlign-70B and GPT-5, meaning that the model’s full utility has already been unlocked when PI security is turned on during the prompt-data channel separation, see Table˜8. However, this is not the case for GPT-4o-mini, GPT-4o, Gemini-2-Flash, Gemini-2.5-Flash, and Gemini-3-Pro, where putting prompt and data in separate message types for PI security noticeably hurts the utility.
Meta SecAlign achieves this good property (no utility drop for security) by using the new input message role for untrusted texts in a free-form manner. For PI security, any texts that were put in the user role can now be directly put within input delimiters if they are untrusted, see Section˜3.
For all tested commercial LLMs, the prompt-data separation is established between the user role and the tool role [wallace2024hierarchy, shi2025lessons], which only treats the tool return as the untrusted data. In our evaluations, we have to create a dummy tool that returns the data texts, and asks the model to call it so that the untrusted data can be processed securely as designed. We cannot know if this design correlates with "the utility drop for PI security" as those models are proprietary, but we hypothesize that the system designs for tool-calling may lead to utility drop in instruction-following tasks. GPT-5 may address this issue by routing proper sub-models to solve corresponding tasks, so we do not observe any utility difference whether PI security is turned on or not.
4.7 Stronger LLMs are more vulnerable to prompt injections if left undefended
We study the relationship between an LLM’s instruction-following capability and its vulnerability to PIs, and investigate the scaling effect of SecAlign++. Intuitively, as LLMs become more capable at instruction following, any injected instructions can be easily identified, leading the model to be more eager to respond to any instruction in its context and thus more vulnerable to PI attacks.
Ideally, we would like to study this by fine-tuning different sizes of LLMs with the same complete (standard) post-training recipe and data. This is not feasible due to our resource constraints. As an alternative, we conduct a proxy study, assuming different sizes of instruction-tuned LLMs within the same model series adopt similar post-training recipe and data. Specifically, we test the undefended Llama-3.1-8B-Instruct, Llama-3.1-70B-Instruct, and Llama-3.3-70B-Instruct, sorted by instruction-following capability in ascending order.
Results in Figure˜4 (left) support our hypothesis that without defense, stronger LLMs suffer from consistently higher ASRs. Fortunately, Figure˜4 (right) shows that after our SecAlign++, all three LLMs can reach a similarly good level of security. This trend gives us hope that, as frontier LLMs continue to improve in capabilities, it is still possible to effectively secure them against PI attacks. However, defenders must move quickly; otherwise, a strong undefended model will become a perfect target for attackers.
5 Discussion
5.1 Conclusion
Model-level defenses are a powerful way to mitigate prompt injection attacks, the top threat to LLM agents. Compared to system-level defenses, they offer strong security and no test-time overhead, as serving a fine-tuned model is as fast as serving the untuned counterpart. Prior model-level defenses are either tested on toy 8B LLMs in academic papers or deployed in closed-source industry APIs (except the recent open-weight-only GPT-OSS [agarwal2025gpt] after our release), but there is no prior open recipe for training a commercial-grade LLM with SoTA security against PIs.
@@ 5.1 Conclusion
Our work bridges this gap by fully open-sourcing a commercial-grade robust foundation LLM, Meta-SecAlign-70B, for secure agentic applications. Our training recipe, SecAlign++, achieves significant security with no noticeable utility drop for the first time in the most comprehensive evaluations to date. Interestingly, this security generalizes to diverse downstream tasks unseen in SecAlign++, especially in agentic workflows where prompt injection is the major threat.
Researchers can apply SecAlign++ to more advanced LLMs than Llama-3.3-70B-Instruct to build more powerful secure foundation models. We hope our work can inspire and accelerate the co-development of PI attacks and defenses in the community.
5.2 Limitations
We focus on defending against (indirect) PIs, where the user is benign, but the environment is malicious, as in agents. Thus, our work cannot prevent jailbreaks [zou2023universal], direct prompt injections [mu2025closer], and other attacks. Despite significant security against static attacks, our model, similar to all existing defenses, is still vulnerable to strong adaptive attacks, see Table˜6 and recent work [nasr2025attacker, wen2025rl, pandya2025may]. That is, the PI threat is far from being solved.
5.3 Future Work
Defending against stronger prompt injections. As stated above, the community still lacks defenses that can resist the current strongest adaptive attacks.
@@ 5.3 Future Work
Online reinforcement learning for securing reasoning LLMs. Our training recipe relies on offline preference optimization to build PI security policy into the non-reasoning model. Recently, online reinforcement learning, such as GRPO [guo2025deepseek], has unlocked better model reasoning using limited data. Reasoning LLMs, with stronger instruction-following abilities, tend to be more vulnerable to PIs [evtimov2025wasp]. Applying online reinforcement learning to securing reasoning LLMs against PI attacks, e.g., by designing PI security rewards, may enjoy similar benefits, with the major challenge lying in utility preservation.
@@
Defending against visual prompt injections. In contrast to textual attacks, the image modality is continuous in nature, and vision models have traditionally been more vulnerable to adversarial manipulation [carlini2019evaluating]. Research towards this problem is especially relevant for agentic web navigation, as SoTA web agents such as OpenAI Operator [openai_operator_system_card], Claude Computer Use [anthropic_claude_computer_use], and Google DeepMind’s Project Mariner [deepmind_project_mariner] are typically powered by multi-modal models.
@@ Acknowledgments
Ethical Considerations
This work develops defense techniques against prompt injection attacks, which is a critical security threat to LLM-integrated applications. Our research aims to improve the security posture of AI systems and does not introduce new attack capabilities. The models and techniques we develop are intended to protect users and systems from malicious manipulation.
We acknowledge that any security research involves dual-use considerations. However, our focus on defense, along with the open-source release of our secure models, is intended to benefit the broader security community by enabling further research into robust defenses. We do not release any novel attack techniques that could be misused.
Our evaluation uses publicly available benchmarks that simulate realistic attack scenarios without causing harm to real systems or users. All experiments were conducted in controlled environments.
@@
In the main paper, we test GPT-5 at its high reasoning level. In Table˜9, we present results at other reasoning levels, which show similar security scores.
Figure˜3 presents an easy utility-security trade-off by tuning LoRA $\alpha$ at test time. This trade-off is traditionally controlled by tuning the learning rate at training time; see Figure˜5.
In the main paper, we put sandwich prompting (Please always remember that your task is: {instruction}) at the end of the data in InjecAgent and AgentDojo. In Table˜10, we present results without the sandwich defense, which also support Meta SecAlign’s advantage over commercial LLMs.
Table 9: GPT-5 with different reasoning levels.
@@
MMLU-Pro ($\uparrow$) - 80.6% 87.1%
GPQA Diamond ($\uparrow$) - 67.3% 85.1%
AlpacaEval2 Utility ($\uparrow$) 67.8% 63.7% 68.7%
@@
Figure 5: Tuning learning rate at training time can also control the utility (bottom) - security (top) trade-off for SecAlign++ on Llama-3.3-70B-Instruct.
Table 10: InjecAgent and AgentDojo results without sandwich prompting defense, which is added to those two benchmarks in the main paper.
Llama-3.3-70B GPT Gemini
Undef. Ours 4o-mini 4o 5 2-Flash 2.5-Flash 3-Pro
InjecAgent ASR ($\downarrow$) 86.0% 2.1% 7.7% 36.9% 0.6% 66.5% 3.5% 2.1%
AgentDojo Utility ($\uparrow$) 62.9% 79.4% 70.1% 80.4% 83.5% 44.3% 58.8% 93.8%
AgentDojo Utility w. Attack ($\uparrow$) 41.9% 77.1% 40.6% 38.8% 81.3% 37.3% 42.9% 89.0%
AgentDojo ASR ($\downarrow$) 23.0% 2.3% 30.9% 43.2% 0.2% 12.4% 30.7% 3.8%
Table 11: Numbers in Figure˜3: Performance (%) using different LoRA $\alpha$ in Meta-SecAlign-70B
@@
Knowledge MMLU ($\uparrow$) 86.3% 86.5% 86.2% 86.1% 86.0% 85.9% 85.9% 85.8% 85.9%
@@
GPQA Diamond ($\uparrow$) 50.0% 48.0% 49.0% 49.0% 47.0% 48.5% 49.0% 47.5% 48.0%
Instruction Following AlpacaEval2 Utility ($\uparrow$) 44.2% 44.4% 45.5% 44.8% 45.0% 44.1% 44.5% 44.6% 44.7%
@@
AlpacaFarm Basic Adaptive ASR ($\downarrow$) 98.1% 97.6% 96.2% 87.0% 59.1% 34.6% 6.3% 1.0% 0.5%
@@
Agentic Workflows InjecAgent ASR ($\downarrow$) 53.6% 39.8% 20.9% 9.5% 4.2% 3.2% 1.3% 0.9% 0.2%
@@
Table 12: Numbers in Figure˜4: security and utility evaluations on LLMs with different capabilities.
Llama-3.1-8B Llama-3.1-70B Llama-3.3-70B
@@
GPQA Diamond ($\uparrow$) 31.3% 28.3% 44.4% 43.9% 50.0% 48.0%
AlpacaEval2 Utility ($\uparrow$) 31.2% 31.0% 43.6% 43.9% 44.2% 44.7%
@@
AlpacaFarm Basic Adaptive ASR ($\downarrow$) 89.4% 0.5% 94.7% 0% 98.1% 0.5%